The ratio of Tom’s age to Mary is 3:4, if Tom is 15, how old is Mary?
Answer by ikleyn(52787)

.
According to the condition,  Tom is 3x years old, while Mary is 4x years old, where x is the common measure of their ages.


Since  3x = 15  (Tom's age), it implies  x= 15/3 = 5.


Hence, Mary is 4*5 = 20 years old.    ANSWER.